### v1

@@ Abstract
Foundation models applied to bio-molecular space hold promise to accelerate drug discovery. Molecular representation is key to building such models. Previous works have typically focused on a single representation or view of the molecules. Here, we develop a multi-view foundation model approach, that integrates molecular views of graph, image and text. Single-view foundation models are each pre-trained on a dataset of up to 200M molecules and then aggregated into combined representations. Our multi-view model is validated on a diverse set of 18 tasks, encompassing ligand-protein binding, molecular solubility, metabolism and toxicity. We show that the multi-view models perform robustly and are able to balance the strengths and weaknesses of specific views. We then apply this model to screen compounds against a large ($>100$ targets) set of G Protein-Coupled receptors (GPCRs). From this library of targets, we identify 33 that are related to Alzheimer’s disease. On this subset, we employ our model to identify strong binders, which are validated through structure-based modeling and identification of key binding motifs.
@@ 1 Introduction
Drug discovery is a complex, multi-stage process. Lead identification and lead optimization remain costly with low success-rates and computational methods play an important role in accelerating these tasks [1, 2, 3]. The prediction of a broad range of chemical and biological properties of candidate molecules is an essential component of screening and assessing molecules and data-driven, machine learning approaches have long aided in this process [4, 5, 6].
@@
Molecular representations form the basis of machine learning models [7, 2], facilitating algorithmic and scientific advances in the field. However, learning useful and generalized latent representation is a hard problem due to limited amounts of labeled data, wide ranges of downstream tasks, vast chemical space, and large heterogeneity in molecular structures. Learning latent representations using unsupervised techniques is vital for such models to scale. Large language models (LLMs) have revolutionized other fields [8] and similar sequence-based foundation models have shown promise to learn molecular representations and be trainable on many downstream property prediction tasks [9, 10, 11]. A key advantage is that the transformer based architecture can learn in a self-supervised fashion to create a “pre-trained” molecular representation. The most direct application of LLM like transformers is facilitated by a sequence, text-based representation (e.g. SMILES).
@@
While sequence-based representations have been hugely successful in capturing human-generated knowledge/information in LLMs, when it comes to representing natural objects like molecules, it’s likely that sequence alone will not be sufficient, and multiple views will be better suited. Here we propose Multi-view Molecular Embedding with Late Fusion (MMELON) a flexible approach to aggregate multiple views in a foundation model setting and show that the resulting richer representations lead to better versatility across a large range of downstream tasks. While in this work, the molecule use case demonstrates the effectivity of the approach, it can be easily extended to include proteins and other biological entities. Our multi-View model is available on GitHub (https://github.com/BiomedSciAI/biomed.multi-view) and Hugging Face (https://Huggingface.co/ibm/biomed.sm.mv-te-84m).
@@
For molecules, the reason sequence alone falls short is that intuitively, molecular representations should reflect the intrinsic geometry of the entities, including symmetries and similarity relations. Such geometry and composition give rise to chemical and biological properties [12]. Indeed, traditional cheminformatics makes use of molecular fingerprints [13, 14, 15, 16, 17], akin to hand-engineered features in other fields.
@@
Graph representations are naturally conducive to molecular symmetries and learn flexible representations [18, 19, 20, 21, 2] However graph neural network (GNN) architectures lack the capacity to scale to large numbers of parameters and learn from very large datasets. Architectures that combine elements of graphs and transformer based approaches are ideal candidates for foundation models. Graph-transformers can represent graphs in a high-capacity neural network and several variants of this approach have been explored [22, 23].
@@
Each representation/modality described above may be considered a ‘view’ of the same molecule which can provide complementary information. A promising approach to achieve more comprehensive molecular representations is identifying effective ways of combining them through multi-modal fusion. Multi-modal fusion is an active area of machine learning across many disciplines. While the molecular use-case has been explored in several publications, primarily in the context of combining graph with a text-based model or 2-dimensional and 3-dimensional molecular representations [26, 27, 28, 29], the scaling of such approach to large scale pre-training ($>100\text{M}$) and the quantification of the value of such combinations on downstream drug discovery tasks remain an open question.
Our multi-view molecular foundation model architecture, MMELON (Fig. 1A), leverages the suite of Biomedical Foundation Model (BMFM) technologies [30]. This multi-view architecture combines molecular representations to learn a base joint multi-modal embedding. This base embedding is then used to fine tune downstream tasks for chemical and biological property prediction. The model currently integrates three views: image, 2-dimensional chemically-bonded graph and text, although our architecture is extensible and can easily include other modalities such as 3-D graphs. For the image view we adopt ImageMol [25], a state of the art molecular foundation model trained on images of molecules. Graph and text views are each pre-trained on a 200M molecular dataset curated from public sources [31, 32]. Using a pre-trained aggregator module to combine the embeddings, we show that the individual modalities contain complimentary information. The weight of each view is a learned parameter of both the pre-trained and fine-tuned models, yielding a transparent interpretation. Our multi-view foundation model is fine-tuned on an extensive set of downstream tasks for molecular property prediction [33, 34, 35]. To validate our model beyond literature benchmarks, it is fine-tuned on binding assay data collected for over 100 G Protein-Coupled receptors (GPCR) [36, 37]. GPCRs associated to Alzheimer’s disease are identified and employed to discover strong binders from a set of metabolites and FDA approved drugs [38]. The predictions are validated by structure-based in silico experiments that identify pharmacophores and binding modes.
@@ 2.2 Pre-trained models learn comprehensive embeddings
Image, Graph and Text are intrinsically different views of chemical structure. However, it is expected that there is at least some overlap in their description of the embedding space. In order to characterize the pre-trained embeddings, we compute the correlation between their Euclidean distances in Fig. 1C. The embeddings are derived from subset of $10,0000$ molecules sampled from the pre-training set. It can be seen that of the three single-view embeddings, Image is the most dissimilar. Text and Graph are highly correlated to each other ($c=0.7$), indicating that these two representations have a similar description of molecular similarity.
@@
Pre-trained molecular embeddings should ideally capture important aspects of the chemical structures of the underlying molecules. While there is no reliable way of directly measuring this, comparison with fingerprints that have been developed by domain experts can serve as a sanity check on whether pre-trained embeddings are in line with chemical knowledge  [11]. In Fig. 1D the correlation coefficients of the single-view models with four common fingerprint descriptors (Morgan, atom pair, MACCS and topological torsion) are plotted. [40, 13, 14, 15]. As a point of reference, we first compute the correlation between these fingerprints, showing that they are, in general, only moderately correlated and yield a different description of the molecules (see Fig 1C). The atom pair fingerprint exhibits the highest correlation to the others.
@@
The correlations plotted in Fig. 1D are computed between the Euclidean distance of the pre-trained models and Tanimoto distance of the fingerprints over the same molecules used above. It can be seen that there is moderate correlation between pre-trained embedding and the fingerprints, demonstrating the validity of the embeddings. Overall, Text exhibits the highest correlation, followed by Graph and Image. Image is most correlated to the MACCS fingerprint, which is expected given that it is used in one of the ImageMol model’s pre-training tasks [25]. Finally, we note that the correlation with fingerprints only serves as a sanity check and strong agreement with fingerprints is not necessarily desirable for all down stream property prediction tasks, because each fingerprint is tailored using a limited set of hand-engineered features. A key objective of learning pre-trained multi-view embeddings is to achieve richer, more flexible representations that are tunable across many downstream tasks.
@@
Figure 1: (A) Schematic of multi-view architecture. Embeddings from three single view pre-trained encoders (Image, Graph and Text) are combined by the aggregator module into a combined embedding. The network is fine-tunable for downstream predictive tasks. (B) Pre-training strategies for single view (Image, Text (sequence) and Graph) embeddings. These strategies are described in detail in Sec. 4.4. (C) The elements of the correlation matrix between the three-single view embeddings (lower triangle) and between four fingerprints (upper triangle). (D) The correlation between four fingerprints and the pre-trained embeddings.
@@ 2.3 Baseline performance: fine-tuning on downstream tasks
The prediction of electronic properties, ligand-protein binding affinity, pharmokinetic properties, and toxicity are examples of tasks that are essential to chemistry and drug discovery. MoleculeNet [34] is a collection of datasets covering such tasks that serves as a common test for machine learning models (see Table S1). Performance on this suite of benchmarks is neither decisive, nor necessarily reflective of all important use-cases in computational drug discovery, but currently it stands as the most widely used point of comparison between models. In addition to MoleculeNet tasks, we assess our model on the task of predicting the inhibition of cytochrome P-450 isoforms [33], which are important in drug metabolism pathways and identifying drug-drug interactions.
In addition to the choice of dataset, the choice of data splitting is critical in robust assessment of model performance. The size-ordered scaffold split (referred to here as ‘scaffold’), clusters molecules according to their Murko scaffold [41] and assigns the clusters to splits in size-order. This is to be compared with other variations of scaffold split, which cluster the dataset in the same manner but assign clusters to train, validation, and test partitions randomly or ‘balanced’ (see, e.g. Ref.  [25]). Scaffold splits are generally understood to be more appropriate and harder than random splits, though the latter are still often utilized in benchmarks [34].
@@
In Fig. 2A, we show the performance of our multi-view model on a wide selection of MoleculeNet and other tasks. The scaffold split is used in all experiments. In order to evaluate the contributions from each view/modality, we focus on comparisons of single view Graph, Text, and Image models and the multi-view model. All are fine-tuned in the same code base with hyper-parameter optimization (primarily the learning rate, see Sec. 4.8). The area under the receiver-operator curve (ROC-AUC) is the chosen metric for the classification tasks and either the root mean square error (RMSE) or mean absolute error (MAE) is chosen for regression as prescribed in Ref. [34]. Regression metrics (Freesolv, Lipophilicity, QM7 and ESOL) are scaled in Fig. 2A so as to be plotted alongside ROC-AUCs (see Sec. 4.8). Tables S2-S3 report the numerical values for experiments with $95\%$ confidence intervals, alongside results trained using a graph convolutional network (GCN) implemented with PyTorch Geometric [42].
@@
We find that overall the multi-view model performs robustly across all tasks, spanning areas such as molecule solubility prediction (ESOL, Lipophilicity), binding activity assay-based classification (HIV, MUV) and toxicity-based tasks such as Tox21 and ClinTox. Screening molecules according to their metabolic properties is particular important in the drug discovery process. To this end, we fine-tune our model on classification of cytochrome-P450 (CYP) inhibition for five CYP isoforms [33] using the balanced scaffold split. Performance as measured by ROC-AUC is high, ranging from $0.90$ to $0.82$ and results are comparable with SOTA performance [25].
@@
In addition, we consider a task that combines our multi-view embedding of small molecules with a popular sequence-based protein embedding, ESM [43]. It is an open question what the protein embedding contributes to so-called ‘drug-target interaction’ tasks [44, 20, 45]. However as ligand embeddings are known to drive model performance, then this still serves as a test of our small-molecule models. We fine-tuned our multi-view and the single-view models on the Davis dataset [35] describing the binding of 72 kinase inhibitors with 442 kinases. We find that the performance of our model exceeds that of a SOTA GNN implementation [45] (Fig. S3).
@@
One advantage of the multi-view late-fusion model architecture is that the contributions of single views to each fine-tune task can be readily recovered through the aggregator weights, $\alpha$ (see Eqn. 3). The $\alpha$ values are visualized as a heat map for MoleculeNet tasks in results in Fig. 2B. We note that the pre-trained aggregator (see Sec. 4.5) is free to evolve during fine-tuning and adapt to the downstream tasks. The Graph model is heavily weighted for classification tasks, which is consistent with the fact that Graph is overall the best performing single-view model for these tasks. In ClinTox, where performance is boosted over Graph (see Fig 2), a more even mixture of embeddings contribute. For regression, the multi-view model draws primarily from Graph and Image views (excepting QM7). A complementary analysis for selected fine-tuned model outputs is given in Sec. 4.7 and Fig. S4.
@@
With few exceptions the Text Model is a minor contributor to the multi-view fine-tuned embeddings. This could be due to it being rather similar to Graph (see Sec. 2.2) and the Graph model is a stronger single view model in our hands. This observation requires further study and could be related to previous results shown for multi-modal architectures [28]. Overall, it can be stated that Graph and Image represent more distinct embeddings (see section 2.2) and are more complimentary.
@@
Figure 2: (A) Performance of Graph (blue line), Image (orange line), Text (green line) and our multi-view model (red line) fine-tuned on diverse downstream tasks. Train, validation, and test splitting is performed with a size-ordered scaffold split. Classification tasks are characterized by ROC-AUC. Regression tasks are characterized by MAE (QM7) and RMSE (others). Regression metrics are scaled so that they can be plotted with ROC-AUC values. (B) Heat map of the weights $\alpha$ that the multi-view model assigns to Graph, Text and Image model for a subset of tasks shown in panel A.

### v3

@@ Abstract
Foundation models applied to bio-molecular space hold promise to accelerate drug discovery. Molecular representation is key to building such models. Previous works have typically focused on a single representation or view of the molecules. Here, we develop a multi-view foundation model approach, that integrates molecular views of graph, image and text. Single-view foundation models are each pre-trained on a dataset of up to 200M molecules and then aggregated into combined representations. Our multi-view model is validated on 22 tasks, encompassing ligand-protein binding, molecular solubility, ADME and toxicity. We show that the multi-view models perform robustly and are able to balance the strengths and weaknesses of specific views. We then apply this model to screen compounds against over$>100$ G Protein-Coupled receptors (GPCRs). We prioritize 33 targets that are related to Alzheimer’s disease and employ our model to identify strong binders. They are validated through structure-based modeling and identification of key binding motifs.
@@ 1 Introduction
Drug discovery is a complex, multi-stage process. Lead identification and lead optimization remain costly with low success-rates [1]. The prediction of a broad range of chemical and biological properties of candidate molecules is an essential component of screening and assessing molecules and data-driven, machine learning approaches have long aided in this process [2, 3, 4, 5, 6].
@@
Molecular representations form the basis of machine learning models [7, 2]. However, learning useful and generalized latent representation is a hard problem due to limited amounts of labeled data, wide ranges of downstream tasks, vast chemical space, and large heterogeneity in molecular structures. Learning latent representations using unsupervised techniques is vital for such models to scale. Large language models (LLMs) have revolutionized many fields [8] and similar sequence-based foundation models have shown promise to learn molecular representations and be trainable on many downstream property prediction tasks [9, 10, 11]. A key advantage is that the transformer based architecture can learn in a self-supervised fashion to create a “pre-trained” molecular representation. The most direct application of LLM like transformers is facilitated by a sequence, text-based representation (e.g. SMILES).
@@
While sequence-based representations have been successful in capturing human-generated knowledge/information in LLMs, when it comes to representing natural objects like molecules, it is likely that sequence molecular representations alone will not be sufficient. Here we propose Multi-view Molecular Embedding with Late Fusion (MMELON) a flexible approach to aggregate multiple views in a foundation model setting and show that the resulting richer representations lead to better versatility across a large range of downstream tasks. The molecule use case demonstrates the effectiveness of the approach, which can be easily extended to include proteins and other biological entities. Our multi-view model is available on GitHub (https://github.com/BiomedSciAI/biomed-multi-view) and Hugging Face (https://Huggingface.co/ibm/biomed.sm.mv-te-84m).
@@
For molecules, the reason sequence alone may fall short is that intuitively, molecular representations should reflect the intrinsic geometry of the entities, including symmetries and similarity relations. Such geometry and composition give rise to chemical and biological properties [12]. Indeed, traditional cheminformatics makes use of molecular fingerprints [13, 14, 15, 16, 17], akin to hand-engineered features in other fields.
@@
Graph representations are naturally conducive to capturing molecular symmetries and other structural properties in a flexible manner [18, 19, 20, 21, 2] However graph neural network (GNN) architectures lack the capacity to scale to large numbers of parameters and learn from very large datasets. Architectures that combine elements of graphs and transformer based approaches are better candidates for foundation models. Graph-transformers can represent graphs in a high-capacity neural network and several variants of this approach have already been explored [22, 23].
Image-recognition is another field where molecular representations can draw from. Image-based representation paired with a convolutional neural network (CNN) architecture have been explored with promising results [24]. ImageMol, an exemplar of this approach, shows strong performance across many downstream tasks [25].
Each representation/modality described above may be considered a ‘view’ of the same molecule which can provide complementary information. A promising approach to achieve more comprehensive molecular representations is identifying effective ways of combining these views through multi-modal fusion. Multi-modal fusion is an active area of machine learning across many disciplines. While the molecular use-case has been explored in several publications, primarily in the context of combining graph with a text-based model or 2-dimensional and 3-dimensional molecular representations [26, 27, 28, 29], the scaling to large pre-training datasets ($>100\text{M}$) and the quantification of the value of such combinations on downstream drug discovery tasks remain open to date.
Our multi-view molecular foundation model architecture, MMELON (Fig. 1A), leverages the suite of Biomedical Foundation Model (BMFM) technologies [30]. This multi-view architecture combines molecular representations to learn a joint base multi-modal embedding. This base embedding is then used to fine tune downstream tasks for chemical and biological property prediction. The model currently integrates three views: image, 2-dimensional chemically-bonded graph and text, although our architecture is extensible and can easily include other modalities such as 3-D conformations. For the image view we adopt ImageMol [25], a state of the art molecular foundation model trained on images of molecules. Graph and text views are each pre-trained on a 200M molecular dataset curated from public sources [31, 32]. Using a pre-trained aggregator module to combine the embeddings, we show that the individual modalities contain complimentary information. The weight of each view is a learned parameter, yielding a transparent interpretation. Our multi-view foundation model is fine-tuned on an extensive set of downstream tasks for molecular property prediction [33, 34, 35].
@@ 2.2 Pre-trained models learn comprehensive embeddings
Image, Graph and Text offer intrinsically different views of chemical structure. However, it is expected that there is at least some overlap in their description of the embedding space. In order to characterize the pre-trained embeddings, we compute the correlation between their Euclidean distances in Fig. 1C. The embeddings are derived from a subset of $10,0000$ molecules sampled from the pre-training set. It can be seen that of the three single-view embeddings, Image is the most distinct. Text and Graph are highly correlated to each other ($c=0.7$), indicating that these two representations have a similar description of molecular similarity.
@@
Pre-trained molecular embeddings should ideally capture important aspects of the chemical structures of the underlying molecules. While there is no absolute way of measuring this, comparison with fingerprints that have been developed by domain experts can serve as a sanity check on whether pre-trained embeddings are in line with chemical knowledge  [11]. In Fig. 1D the correlation coefficients of the single-view models with four common fingerprint descriptors (Morgan, atom pair, MACCS and topological torsion) are plotted. [48, 13, 14, 15]. As a point of reference, we first compute the correlation between these fingerprints, showing that they are moderately correlated and yield a different description of the molecules (see Fig 1C). The atom pair fingerprint exhibits the highest correlation to the others.
@@
The correlations plotted in Fig. 1D are computed between the Euclidean distance of the pre-trained models and Tanimoto distance of the fingerprints over the same molecules used above. It can be seen that there is moderate correlation between pre-trained embedding and the fingerprints, demonstrating the validity of the embeddings. Overall, Text exhibits the highest correlation, followed by Graph and Image. Image is most correlated to the MACCS fingerprint, which is expected given that it is used in one of the ImageMol model’s pre-training tasks [25]. Finally, we note that the correlation with fingerprints only serves as a sanity check and strong agreement with fingerprints is not necessarily desirable for all down stream property prediction tasks, because each fingerprint is tailored using a limited set of hand-engineered features. A key objective of learning pre-trained multi-view embeddings is to achieve richer, more flexible representations.
@@
Figure 1: (A) Schematic of multi-view architecture. Embeddings from three single view pre-trained encoders (Image, Graph and Text) are combined by the aggregator module into a combined embedding. The network is fine-tunable for downstream predictive tasks. (B) Pre-training strategies for single view (Image, Text (sequence) and Graph) embeddings. These strategies are described in detail in Sec. 4.4. (C) The elements of the correlation matrix between the three-single view embeddings (lower triangle) and between four fingerprints (upper triangle). (D) The correlation between four fingerprints and the pre-trained, single-view embeddings.
@@ 2.3 Baseline performance: fine-tuning on downstream tasks
The prediction of electronic properties, ligand-protein binding affinity, pharmokinetic properties, and toxicity are examples of tasks that are essential to chemistry and drug discovery. MoleculeNet [34] is a collection of datasets covering such tasks that serves as a common test for machine learning models. Performance on this suite of benchmarks is neither decisive, nor necessarily reflective of all important use-cases in computational drug discovery, but currently it stands as the most widely used point of comparison between models. We also assess our model on predicting the inhibition of cytochrome P-450 isoforms [33](CYP) which are critical in drug metabolism pathways, and the recently released ComputationalADME dataset, which include tasks such as liver microsomal stability and plasma protein binding [49]. In addition to the choice of dataset, the choice of data splitting is critical in robust assessment of model performance. We primarily adopt scaffold-based methodologies, but also utilize random splits where appropriate (see SI for further discussion).
@@
In Fig. 2A, we show the performance of our multi-view model on 21 tasks from MoleculeNet, CYP and ComputationalADME. In order to evaluate the contributions from each view/modality, we focus on comparisons of single view Graph, Text, and Image models and the multi-view model. To serve as a simple baseline, results from a graph convolutional network (GCN) implemented with PyTorch Geometric are shown [50]. All are fine-tuned in the same code base with hyper-parameter optimization (primarily the learning rate, see Sec. 4.8). The area under the receiver-operator curve (ROC-AUC) is the chosen metric for the classification tasks and root mean square error (RMSE) or mean absolute error (MAE) is chosen for regression [34]. Regression metrics are scaled in Fig. 2A so as to be plotted alongside ROC-AUCs (see Sec. 4.8). Tables S2-S5 report the numerical values for experiments with $95\%$ confidence intervals. A detailed description of each task is given in Table S1.
@@
We find that overall the multi-view model performs robustly across all tasks, spanning areas such as molecule solubility prediction (ESOL, Lipophilicity), binding activity assay-based classification (HIV, MUV) and ADME-Tox based tasks such as Tox21 and the ComputationalADME set (e.g. RPPB, HPPB, RLM, HLM). In particular, we test on cytochrome-P450 (CYP) inhibition for five CYP isoforms [33]. Performance as measured by ROC-AUC is high, ranging from $0.90$ to $0.82$ and results are comparable with SOTA performance [25]. For example, isoform CYP2C9 is one of most important contributor to drug metabolism [51] and our model achieved a high ROC-AUC value (0.90, Fig. 2A). The binding modes of top bio-active molecules computed from molecular docking align well with reported structures of CYP2C9 in complex with the drug losartan [52] (PDB ID: 8VX0, see Fig. 2B). In addition, the salient molecular moieties as predicted by the graph and image views of multi-view are both close to the catalytic site of CYP2C9 [53], suggesting our multi-view models could highlight the important binding features.
@@
In addition, we consider a task that combines our multi-view embedding of small molecules with a popular sequence-based protein embedding, ESM [54]. It is an open question what the protein embedding contributes to so-called ‘drug-target interaction’ tasks [55, 20, 56]. However, as ligand embeddings are known to drive model performance, this could still serve as another test of our small-molecule models. We fine-tuned our multi-view and the single-view models on the Davis dataset [35] describing the binding of 72 kinase inhibitors with 442 kinases. We find that the performance of our model exceeds that of a SOTA GNN implementation [56] (Fig. S3).
@@
One advantage of the multi-view late-fusion model architecture is that the contributions of single views to each task can be readily recovered through the aggregator weights, $\alpha$ (see Eqn. 3). The $\alpha$ values are visualized as a heat map for MoleculeNet tasks in results in Fig. 2C. We note that the pre-trained aggregator (see Sec. 4.5) is free to evolve during fine-tuning and adapt to the downstream tasks. The Graph model is heavily weighted for most tasks, which is consistent with the fact that Graph is overall the best performing single-view model. A complementary analysis measuring the agreement between different views of selected fine-tuned tasks is given in Sec. 4.7 and Figs. S4-S5. For CYP models, both Graph and Image are nearly equally correlated with the multi-view model, and this degeneracy can be related to trends in Fig. 2C.
@@
With the exception of QM7, the Text Model is a minor contributor to the multi-view fine-tuned embeddings and while the predictions between text and multi-view are correlated, they exhibits the weakest relation (see Figs. S4-S5). This could be due to it being rather similar to Graph (see Sec. 2.2) while the Graph model is a stronger single view model. This observation requires further study and could be related to previous results shown for multi-modal architectures [28]. Overall, it can be stated that Graph and Image represent more distinct embeddings (see section 2.2) and are more complimentary.
@@
Figure 2: (A) Performance of GCN (blue line), Graph (orange line), Image (green line), Text (red line) and multi-view model (purple line) fine-tuned on diverse downstream tasks. Train, validation, and test splitting is performed as described in the text. Classification tasks are characterized by ROC-AUC. Regression metrics (RMSE, MAE) are scaled so that they can be plotted with ROC-AUC values. (B) Docking poses of selected bioactive molecules against CYP2C9 as predicted by the multi-view model are compared with the reported crystal structure (PDB: 8VX0, left column panels) and are shown alongside graph (center column panels) and image (right column panels) attention level heatmaps. Darker color indicates higher attention level. (C) Heat map of the weights $\alpha$ that the multi-view model assigns to Graph, Text and Image model for a subset of tasks shown in panel A.